OpenStudy (anonymous):

If cos Θ = negative two over five and tan Θ > 0, what is the value of sin Θ? negative square root 21 negative square root 21 over 5 square root 21 over 5 square root of 21 over 2

OpenStudy (anonymous):



OpenStudy (anonymous):

\[a^2 + b^2 = c^2\]\[(-2)^2 + x^2 = 5^2\]\[x^2 = 25 - 4\]\[x = \sqrt{21}\] \[\sin(Θ) = x/5\]\[\sin(Θ) = -\sqrt{21}/5\]

OpenStudy (anonymous):

So, negative square root 21 over 5

OpenStudy (anonymous):

so you just plugged the into the pythagoren theorem, I get it now :) ty

OpenStudy (jdoe0001):

\(\bf cos(\theta)=-\cfrac{2}{5}\qquad tan(\theta)>0 \\ \quad \\ tan(\theta)>0\Leftarrow\textit{another way of saying, tangent is positive}\\ \textit{meaning, 3rd or 1st Quadrants} \\ \quad \\ cos(\theta)=-\cfrac{2}{5}\implies \cfrac{adjacent}{hypotenuse}\implies \cfrac{a=-2}{c=5}\) notice ASNinja 's picture coordinates